# A battle of designs: triangular vs. L-shaped detectors and parity violation in the gravitational-wave background

### v1

@@ Abstract
We investigate the prospects for detecting a parity-violating gravitational-wave background (GWB) with third-generation ground-based detector networks. We focus on a network consisting of one Einstein Telescope (ET) and two Cosmic Explorer (CE) detectors. In our analysis we vary the ET design, detector orientations, and arm lengths, in order to assess the impact of geometry and scale on detection capabilities. We demonstrate that networks with an L-shaped ET design have stronger parity violation constraining power than networks with a triangular ET design, particularly seen when studying ET designs on their own.
@@ I Introduction
The superposition of independent gravitational waves (GW) of cosmological and astrophysical origin forms a gravitational-wave background (GWB). The detection of such a background can shed light on early-universe mechanisms, such as inflation Grishchuk (1974); Starobinsky (1979); Badger et al. (2024), first-order phase transitions Kamionkowski et al. (1994a); Apreda et al. (2002); Grojean and Servant (2007); Romero et al. (2021); Badger et al. (2023), and cosmic strings Auclair et al. (2020); Damour and Vilenkin (2001); Abbott et al. (2021a), and aid in characterizing astrophysical compact binary coalescences (CBCs) sources and populations Abbott et al. (2023). Searches for an unpolarized GWB using data from the ground-based interferometers LIGO Aasi et al. (2015) and Virgo Acernese et al. (2015) have placed upper limits on its energy density Abac et al. (2025a).
@@
Probing for polarization in the GWB can place constraints on parity-violating theories Crowder et al. (2013). Parity-violating effects in the GWB have been studied in the context of current and future detector sensitivities Seto (2007); Ricciardone (2017); Orlando et al. (2021); Bartolo et al. (2018); Domcke et al. (2020); Xu et al. (2019); Martinovic et al. (2021a); Badger and Sakellariadou (2021); Cruz et al. (2024), and have been constrained through analyses of LIGO–Virgo data Crowder et al. (2013); Martinovic et al. (2021a); Abac et al. (2025b).
@@ IV Future ground-based detector configurations
We consider seven representative networks, each combining one Einstein Telescope (ET) with two Cosmic Explorer (CE) detectors, using the geometric information from  Ebersold et al. (2024). We include triangular ET designs with 10 and 15 km arms ($\Delta_{10}$, $\Delta_{15}$), as well as ET designs consisting of two L-shaped interferometers with 15 km arms (S, R). For CE, we consider L-shaped detectors with 20 km (P) and 40 km (Y) arms. The triangular configurations consist of three interferometers, while an L-shaped detector consists of one interferometer. Thus a $\Delta$ + (PY) network contains five detectors (10 baselines), while an (SR) + (PY) network contains four (6 baselines). These configurations explore how detector orientations and arm length influence network performance, with some optimized for GWB sensitivity and others for CBC source localization, clarifying whether differences between triangular and L-shaped designs arise from geometry or arm length. Further technical details are in Ebersold et al. (2024), for which we use the amplitude spectral density curves found in Danilishin and Zhang (2023) and Fulda et al. (2023), and the full set of configuration names and descriptions is in Table 1.
@@
In Fig. 1, we present the $I$- and $V$-mode PI curves for the various network configurations. The $\Omega_{{\rm{PI}},I}$ curves show that GWB-optimized networks yield the highest $I$-mode sensitivity, which is expected by construction. In particular, (SR)1(PY)1 dominates below $\sim$ 40 Hz, while at higher frequencies the 15 km triangular configurations, $\Delta_{15}$(PY)4 and $\Delta_{15}$(PY)5, outperform the 2L designs.
The $\Omega_{{\rm{PI}},V}$ curves exhibit reduced sensitivity compared to their $I$-mode counterparts, with sensitivity decreasing as $|\Pi|$ approaches 0. Their ranking is relatively stable across frequencies: below 20 Hz, (SR)3(PY)3 is most sensitive and $\Delta_{10}$(PY)5 least, while above 25 Hz the best sensitivity shifts to (SR)1(PY)1, with $\Delta_{10}$(PY)5 remaining the weakest.
This result is expected since the co-planar triangular ET detectors are weak to $V$-modes Domcke et al. (2020); Badger and Sakellariadou (2021); Seto and Taruya (2008) and weaken a network’s overall sensitivity to PV detection in comparison to networks with 2L ET designs that are spatially separated. This was verified with a separate study in which the PV detection prospects of ET alone were examined, and find that present constraints Abac et al. (2025b) have ruled out the parameter space for which ET detector can detect circular polarization, thus the inclusion of CE detectors is a necessity to make proper parity-violation inferences. See Appendix A for further details.
Figure 1: PI sensitivity curves for II-mode (solid lines) and $V$-mode (dotted and dashed lines), shown for the seven different detector third-generation network configurations for an observing time of one year and $\rho=4$. The $V$-mode curves correspond to $|\Pi|=1$ (dotted) and $|\Pi|=0.1$ (dashed).
@@ V Future detection prospects
We present theoretical (based on signal-to-noise ratio) and data-driven (obtained with a Bayesian analysis) prospects for detecting PV by applying the previously described formalisms to third-generation detector networks. Our analysis covers general power-law GWBs with $12.5\leq\log_{10}\Omega_{\alpha}\leq-10$, spectral indices $\alpha=-3,\dots,3$, and polarization degrees $1\leq\Pi\leq 1$. We note that the upper bound on the $\log_{10}\Omega_{\alpha}$ range is determined by current detector constraints Abac et al. (2025a), while the lower bound is to allow studied GWBs to be potentially detected in the considered networks, as seen in Fig. 1.
@@ V.1 Signal-to-noise ratio
The sensitivities of the third-generation detector networks are quantified by the $I$- and $V$-mode SNRs. Figure 2 shows the results for a flat PL ($\alpha=0$) spectrum across the seven network configurations. As expected, $\rho_{I}$ is independent of $\Pi$ and scales only with the GWB amplitude, while $\rho_{V}$ grows with $|\Pi|$, enhancing detectability for more strongly polarized backgrounds. We repeat this study for other power-law indices and rank their detection prospects in Tables 3 and 3.
Figure 2: SNRs for II-mode (solid lines) and $V$-mode (dashed lines) for the seven third-generation detector network configurations, assuming a flat ($\alpha=0$) power-law GWB.
$\alpha$ 3,2-3,-2 1,0-1,0 $\tfrac{1}{2},\tfrac{2}{3},1,2$ 33
1st (SR)1(PY)1 (SR)1(PY)1 (SR)1(PY)1 $\Delta_{15}$(PY)4
2nd (SR)3(PY)3 $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)5
3rd $\Delta_{15}$(PY)4 (SR)3(PY)3 $\Delta_{10}$(PY)4 $\Delta_{10}$(PY)4
4th $\Delta_{15}$(PY)5 $\Delta_{10}$(PY)4 (SR)3(PY)3 $\Delta_{10}$(PY)5
5th $\Delta_{10}$(PY)4 $\Delta_{15}$(PY)5 $\Delta_{15}$(PY)5 (SR)1(PY)1
6th (SR)2(PY)2 (SR)2(PY)2 $\Delta_{10}$(PY)5 (SR)3(PY)3
7th $\Delta_{10}$(PY)5 $\Delta_{10}$(PY)5 (SR)2(PY)2 (SR)2(PY)2
Table 2: Ranking from best (1st) to worst (7th) of detector networks by II-mode SNRs for different PL indices $\alpha$.
$\alpha$ 3-3 to 11 $2,3$
1st (SR)3(PY)3 (SR)1(PY)1
2nd (SR)1(PY)1 (SR)3(PY)3
3rd $\Delta_{15}$(PY)4 (SR)2(PY)2
4th (SR)2(PY)2 $\Delta_{15}$(PY)4
5th $\Delta_{15}$(PY)5 $\Delta_{15}$(PY)5
@@
We find that the 2L network (SR)1(PY)1 provides the largest $\rho_{I}$ for $3\leq\alpha\leq 2$, while the 15 km triangular network $\Delta_{15}$(PY)4 overtakes it at $\alpha=3$, highlighting the advantage of a triangular design for GWBs with large positive spectral indices. For the $V$-mode, (SR)3(PY)3 is most sensitive for $3\leq\alpha\leq 1$, with (SR)1(PY)1 dominating at larger indices. Among triangular ETs, longer arms ($\Delta_{15}$(PY)4,5) outperform their 10 km counterparts, confirming that arm length impacts sensitivity. Nevertheless, networks incorporating a 2L ET design consistently achieve the highest $\rho_{V}$, as these networks have non-vanishing $\gamma_{V}^{d_{1}d_{2}}$, which significantly enhance sensitivity to the $V$-mode. To verify that this enhancement is intrinsic to the ET geometry rather than the CE orientation, Appendix A presents ET-only configurations. These results confirm that triangular ETs suffer a dramatic loss of PV sensitivity compared to 2L designs. Hence, optimal PV detection requires an ET with a 2L design, highlighting that detector geometry is the dominant factor for detecting parity-violating signals. These results are consistent with the PI curves, whose frequency-dependent shapes qualitatively explain the SNR rankings across different spectral indices. We highlight that fixing different reference frequencies result in the same ordering of SNR prospects, emphasizing the generality of our results.
@@ V.2 Bayesian analysis
We now complement the SNR analysis with a data-driven Bayesian analysis on simulated signals, assuming the design sensitivity and noise characteristics of each third-generation network configuration. Parameter estimation is carried out with the pygwb Renzini et al. (2024) and Bilby  Ashton et al. (2019) packages using the Dynesty sampler. This allows us to assess a network’s capability to detect a polarized signal and to distinguish it from an unpolarized one. To this end, we fix a power-law index $\alpha$ and generate 1600 simulated injections uniformly over the $\log_{10}\Omega_{\alpha}$ and $\Pi$ range in each network configuration, then searching over the priors of Table 4 based off of theoretical bounds and previously explored search priors Martinovic et al. (2021b); Abac et al. (2025b). The procedure is then repeated with $\Pi=0$ to model unpolarized backgrounds, using the same amplitude and spectral index priors. To reduce statistical noise, we generate 15 independent injection sets with random detector noise and combine results with an optimized weighted average Romano and Cornish (2017). This enables a direct comparison between polarized and unpolarized searches.
@@
Table 4: Priors used in the Bayesian analysis. The $\log_{10}\Omega_{\rm\alpha}$ and $\alpha$ priors follow previous searches, while taking into account the sensitivity of third-generation detectors Abbott et al. (2021b); Martinovic et al. (2021b). The prior on $\Pi$ is the full, physically allowed range of values Martinovic et al. (2021a).
@@ V.2.1 Comparison to signal-to-noise
We compare the recovered data-analysis Bayes factors to the $I$- and $V$-mode SNRs for all injections. Figure 3 shows this comparison for the most optimistic ((SR)3(PY)3) and most pessimistic ($\Delta_{10}$(PY)5) network configurations for a flat GWB spectrum. More concretely, it shows the SNRs alongside the log Bayes factors for a polarized model compared to detector noise, $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$, and to an unpolarized model, $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$. We adopt strong detection thresholds of $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$ and $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$ equal to 8. These correspond to $\rho_{I}$ and $\rho_{V}$ equal to 4, since $\rho^{2}\simeq 2\ln\mathcal{B}$ in the strong signal regime Romano and Cornish (2017). Our findings confirm this relation.
@@
In conclusion, the SNR predictions reproduce the trends seen in the data-driven analysis across both network extremes, demonstrating that simple theoretical estimates based on SNRs provide an excellent guide to the detectability of polarized GWBs.
@@ V.2.2 Parameter estimation
We perform a Bayesian analysis to assess parameter recovery of simulated parity-violating GWBs for the most optimistic, (SR)3(PY)3 and most pessimistic, $\Delta_{10}$(PY)5, network configurations. From the marginalized posteriors, we compute the 95% Upper Limits (UL) and Lower Limits (LL) on $\Pi$, allowing us to make $\Pi=0$ exclusion analyses. This is shown in Fig. 3: purple and green regions denote where $\Pi=0$ is ruled out with at least $95\%$ confidence by the 2L and triangular networks, respectively. We define a threshold region as the parameter space where the 2L network excludes $\Pi=0$ at $95\%$ confidence, but the triangle network does not. An average V-mode SNR corresponding to the borders of this threshold region is $\rho_{V}^{\Delta_{10}\text{(PY)}_{5}}\leq 2.16\leq\rho_{V}^{\text{(SR)}_{3}\text{(PY)}_{3}}$, which is indicated by the gray-shaded region in Fig.  3.
@@
Next, we focus on a flat parity-violating GWB case study with $\log_{10}\Omega_{\alpha}=-11.54$, $\Pi=-1$, indicated by the red star in Fig. 3. These parameters were chosen so that the signal lies in the threshold regime. The combined posteriors results are shown in Fig. 4, where model parameters are recovered within $1\sigma$ for both networks. We derive the $95\%$ UL on $\Pi$ equal to 0.238-0.238 and $0.233$ for the 2L and triangular network, respectively. We also compute log Bayes factors for the parity-violating and the unpolarized GWB searches. For the (SR)3(PY)3 network, we obtain $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=2.914\pm 0.043$, while for $\Delta_{10}$(PY)5 we find $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=0.946\pm 0.037$, highlighting the greater capability of (SR)3(PY)3 to distinguish a polarized from unpolarized GWBs than $\Delta_{10}$(PY)5. Although these log Bayes factors do not indicate a strong preference for a PV model, this injection demonstrates that (SR)3(PY)3 can rule out $\Pi=0$ at the $95\%$ confidence level, while $\Delta_{10}$(PY)5 cannot.
In conclusion, these findings highlight that network design strongly impacts parameter constraints, particularly for circular polarization $\Pi$, with the 2L network providing tighter constraints on parity-violating GWBs.

### v2

@@ Abstract
We investigate the prospects for detecting a parity-violating gravitational-wave background (GWB) with third-generation ground-based detector networks. We focus on a network consisting of one Einstein Telescope (ET) and two Cosmic Explorer (CE) detectors. In our analysis we vary the ET design, detector orientations, and arm lengths to assess the impact of geometry and scale on detection capabilities. We find that, among the configurations studied, networks with a $15$ km triangular ET design provide the strongest parity-violation constraining power, followed by networks with an L-shaped ET design, while networks with $10$ km triangular ET configurations are the least sensitive. Under current bounds, ET alone cannot confidently detect parity violation.
@@ I Introduction
The superposition of independent gravitational waves (GW) of cosmological and astrophysical origin forms a gravitational-wave background (GWB). The detection of such a background can shed light on early-universe mechanisms, such as inflation Grishchuk (1974); Starobinsky (1979); Badger et al. (2024), first-order phase transitions Kamionkowski et al. (1994a); Apreda et al. (2002); Grojean and Servant (2007); Romero et al. (2021); Badger and others (2023), and cosmic strings Auclair and others (2020); Damour and Vilenkin (2001); Abbott and others (2021), and aid in characterizing astrophysical compact binary coalescences (CBCs) sources and populations Abbott and others (2023). Searches for an unpolarized GWB using data from the ground-based interferometers LIGO Aasi and others (2015) and Virgo Acernese and others (2015) have placed upper limits on its energy density Abac and others (2025b).
@@
Probing for polarization in the GWB can place constraints on parity-violating theories Crowder et al. (2013). Parity-violating effects in the GWB have been studied in the context of current and future detector sensitivities Seto (2007); Ricciardone (2017); Orlando et al. (2021); Bartolo et al. (2018); Domcke et al. (2020); Xu et al. (2019); Martinovic et al. (2021a); Badger and Sakellariadou (2021); Cruz et al. (2024), and have been constrained through analyses of LIGO–Virgo data Crowder et al. (2013); Martinovic et al. (2021a); Abac and others (2025a).
@@ IV Future ground-based detector configurations
We consider seven representative networks, each combining one Einstein Telescope (ET) with two Cosmic Explorer (CE) detectors, using the geometric information from  Ebersold et al. (2024). We include triangular ET designs with $10$ and $15$ km arms ($\Delta_{10}$, $\Delta_{15}$), as well as ET designs consisting of two L-shaped detectors with $15$ km arms (S, R), following the standard ET configurations introduced in  Branchesi and others (2023) and extended to ET+CE networks in  Ebersold et al. (2024). For CE, we consider L-shaped detectors with $20$ km (P) and $40$ km (Y) arms. The triangular configurations consist of three detectors, while an L-shaped configuration consists of one. Thus a $\Delta$ + (PY) network contains five detectors ($10$ baselines), while an (SR) + (PY) network contains four ($6$ baselines). These configurations explore how detector orientations and arm length influence network performance, with some optimized for GWB sensitivity and others for CBC source localization, clarifying whether differences between triangular and L-shaped designs arise from geometry or arm length. Further technical details are in Ebersold et al. (2024), for which we use the amplitude spectral density curves found in Danilishin and Zhang (2023) and Fulda et al. (2023), and the full set of configuration names and descriptions is in Table 1.
@@
In Fig. 1, we present the $I$- and $V$-mode PI curves for the various network configurations, extending previous ET design studies Branchesi and others (2023) to networks combining ET with CE and to $V$-mode sensitivity. The $\Omega_{{\rm{PI}},I}$ curves show that GWB-optimized networks yield the highest $I$-mode sensitivity, which is expected by construction. In particular, (SR)1(PY)1 dominates below $20$ Hz, while at higher frequencies all triangular configurations outperform the 2L designs.
The $\Omega_{{\rm{PI}},V}$ curves exhibit reduced sensitivity compared to their $I$-mode counterparts, with sensitivity decreasing as $|\Pi|$ approaches 0. Their ranking is relatively stable across frequencies: $\Delta_{15}$(PY)4 is consistently the most sensitive configuration, while $\Delta_{10}$(PY)5 performs the worst. Overall, networks incorporating a $15$ km triangular ET design achieve the highest sensitivity to $V$-modes, followed by 2L ET networks, whereas $10$ km triangular ET configurations are the least sensitive. The superior performance of the $15$ km triangular configuration relative to the $15$ km 2L case arises from its larger number of independent baselines. In the $10$ km triangular configuration, however, higher instrumental noise dominates, leading to reduced sensitivity. Sensitivity to $V$-modes is therefore controlled by a combination of arm length, the number of independent baselines, and network geometry. The parity-violation detection prospects of ET alone are presented in Appendix  A. We find that current constraints Abac and others (2025a) have already excluded the parameter space in which ET alone would be capable of detecting circular polarization. Therefore, the inclusion of CE detectors is a necessity to make robust parity-violation inferences.
Figure 1: PI sensitivity curves for $I$-mode (solid lines) and $V$-mode (dotted and dashed lines), shown for the seven different detector third-generation network configurations for an observing time of one year and $\rho=4$. The $V$-mode curves correspond to $|\Pi|=1$ (dotted) and $|\Pi|=0.1$ (dashed).
@@ V Future detection prospects
We present theoretical (based on signal-to-noise ratio) and data-driven (obtained with a Bayesian analysis) prospects for detecting PV by applying the previously described formalisms to third-generation detector networks. Our analysis covers general power-law GWBs with $12.5\leq\log_{10}\Omega_{\alpha}\leq-10$, spectral indices $\alpha=-3,\dots,3$, and polarization degrees $1\leq\Pi\leq 1$. We note that the upper bound on the $\log_{10}\Omega_{\alpha}$ range is determined by current detector constraints Abac and others (2025b), while the lower bound is to allow studied GWBs to be potentially detected in the considered networks, as seen in Fig. 1.
@@ V.1 Signal-to-noise ratio
The sensitivities of the third-generation detector networks are quantified by the $I$- and $V$-mode SNRs. Fig. 2 shows the results for a flat PL ($\alpha=0$) spectrum across the seven network configurations. As expected, $\rho_{I}$ is independent of $\Pi$ and scales only with the GWB amplitude, while $\rho_{V}$ grows with $|\Pi|$, enhancing detectability for more strongly polarized backgrounds. We repeat this study for other power-law indices and rank their detection prospects in Tables 3 and 3.
Figure 2: SNRs for $I$-mode (solid lines) and $V$-mode (dashed lines) for the seven third-generation detector network configurations, assuming a flat ($\alpha=0$) power-law GWB.
Order $\alpha$ 3,2-3,-2 $1,0,\tfrac{1}{3},\tfrac{1}{2},\tfrac{2}{3},1$ 22 33
1st (SR)1(PY)1 (SR)1(PY)1 $\Delta_{15}$(PY)4
$\Delta_{15}$(PY)4
$\Delta_{15}$(PY)5
2nd $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)4 (SR)1(PY)1
$\Delta_{10}$(PY)4
$\Delta_{10}$(PY)5
3rd $\Delta_{15}$(PY)5 (SR)3(PY)3 $\Delta_{15}$(PY)5 (SR)1(PY)1
4th (SR)3(PY)3 $\Delta_{10}$(PY)4 $\Delta_{10}$(PY)4 (SR)3(PY)3
5th $\Delta_{10}$(PY)4 $\Delta_{15}$(PY)5 (SR)3(PY)3 (SR)2(PY)2
6th (SR)2(PY)2 $\Delta_{10}$(PY)5 $\Delta_{10}$(PY)5
7th $\Delta_{10}$(PY)5 (SR)2(PY)2 (SR)2(PY)2
Table 2: Ranking from best (1st) to worst (7th) of detector networks by II-mode SNRs for different PL indices $\alpha$, grouping together indices that yield identical prospects.
Order $\alpha$ 3-3 to 11 $2,3$
1st $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)4
2nd $\Delta_{15}$(PY)5 $\Delta_{15}$(PY)5
3rd (SR)3(PY)3 (SR)1(PY)1
4th (SR)1(PY)1 (SR)3(PY)3
5th (SR)2(PY)2 (SR)2(PY)2
@@
We find that the 2L network (SR)1(PY)1 provides the largest $\rho_{I}$ for $3\leq\alpha\leq 1$, while the $15$ km triangular network $\Delta_{15}$(PY)4 overtakes it at $\alpha=2$, highlighting the advantage of a triangular design for GWBs with large positive spectral indices. For the $V$-mode, $\Delta_{15}$(PY)4 is the most sensitive configuration across all $\alpha$ values, with the 2L networks consistently outperforming the $10$ km triangular configurations. This hierarchy indicates that optimal parity-violation detection is achieved with a $15$ km triangular ET combined with two CEs. This behavior is consistent with the PI curves in Fig. 1, whose frequency-dependent shapes qualitatively explain the SNR rankings across different spectral indices. We highlight that fixing different reference frequencies yields the same ordering of SNR prospects, emphasizing the generality of our results. Appendix A presents ET-only configurations and shows that all triangular ET designs suffer a dramatic loss of PV sensitivity compared to 2L designs.
@@ V.2 Bayesian analysis
We now complement the SNR analysis with a data-driven Bayesian analysis on simulated signals, assuming the design sensitivity and noise characteristics of each third-generation network configuration. Parameter estimation is carried out with the pygwb Renzini and others (2024) and Bilby  Ashton and others (2019) packages using the Dynesty sampler. This allows us to assess a network’s capability to detect a polarized signal and to distinguish it from an unpolarized one. To this end, we fix a power-law index $\alpha$ and generate 1600 simulated injections uniformly over the $\log_{10}\Omega_{\alpha}$ and $\Pi$ range in each network configuration, then searching over the priors of Table 4 based off of theoretical bounds and previously explored search priors Martinovic et al. (2021b); Abac and others (2025a). The procedure is then repeated with $\Pi=0$ to model unpolarized backgrounds, using the same amplitude and spectral index priors. To reduce statistical noise, we generate $15$ independent injection sets with random detector noise and combine results with an optimized weighted average Romano and Cornish (2017). This enables a direct comparison between polarized and unpolarized searches.
@@
Table 4: Priors used in the Bayesian analysis. The $\log_{10}\Omega_{\rm\alpha}$ and $\alpha$ priors follow previous searches, while taking into account the sensitivity of third-generation detectors Abbott et al. (2021); Martinovic et al. (2021b). The prior on $\Pi$ is the full, physically allowed range of values Martinovic et al. (2021a).
@@ V.2.1 Comparison to signal-to-noise
We compare the recovered data-analysis Bayes factors to the $I$- and $V$-mode SNRs for all injections. Fig. 3 shows this comparison for the (SR)3(PY)3 and $\Delta_{10}$(PY)5 configurations for a flat GWB spectrum. More concretely, it shows the SNRs alongside the log Bayes factors for a polarized model compared to detector noise, $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$, and to an unpolarized model, $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$. We adopt strong detection thresholds of $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$ and $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$ equal to 8. These correspond to $\rho_{I}$ and $\rho_{V}$ equal to 4, since $\rho^{2}\simeq 2\ln\mathcal{B}$ in the strong signal regime Romano and Cornish (2017). Our findings confirm this relation.
@@
In conclusion, the SNR predictions reproduce the trends seen in the data-driven analysis across both networks, demonstrating that simple theoretical estimates based on SNRs provide an excellent guide to the detectability of polarized GWBs. We note that the present analysis assumes uncorrelated detector noise. For the ET triangular configuration, correlated Newtonian noise between the co-located detectors is expected to further degrade the sensitivity to GWBs Janssens and others (2024). Since such correlated noise is absent for spatially separated L-shaped detectors, including this effect would reduce the performance of the triangular network and strengthen the preference for 2L configurations found here.
@@ V.2.2 Parameter estimation
We perform a Bayesian analysis to assess parameter recovery of simulated parity-violating GWBs for the (SR)3(PY)3 and $\Delta_{10}$(PY)5 configurations. From the marginalized posteriors, we compute the 95% Upper Limits (UL) and Lower Limits (LL) on $\Pi$, allowing us to make $\Pi=0$ exclusion analyses. This is shown in Fig. 3: purple and green regions denote where $\Pi=0$ is ruled out with at least $95\%$ confidence by the 2L and triangular networks, respectively. We define a threshold region as the parameter space where the 2L network excludes $\Pi=0$ at $95\%$ confidence, but the triangle network does not. An average $V$-mode SNR corresponding to the borders of this threshold region is $\rho_{V}^{\Delta_{10}\text{(PY)}_{5}}\leq 2.16\leq\rho_{V}^{\text{(SR)}_{3}\text{(PY)}_{3}}$, which is indicated by the gray-shaded region in Fig.  3.
@@
Next, we focus on a flat parity-violating GWB case study with $\log_{10}\Omega_{\alpha}=-11.54$, $\Pi=-1$, indicated by the red star in Fig. 3. These parameters were chosen so that the signal lies in the threshold regime. The combined posterior results are shown in Fig. 4, where model parameters are recovered within $1\sigma$ for both networks. We derive the $95\%$ UL on $\Pi$ equal to 0.238-0.238 and $0.233$ for the 2L and triangular network, respectively. We also compute log Bayes factors for the parity-violating and the unpolarized GWB searches. For the (SR)3(PY)3 network, we obtain $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=2.914\pm 0.043$, while for $\Delta_{10}$(PY)5 we find $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=0.946\pm 0.037$, highlighting the greater capability of (SR)3(PY)3 to distinguish a polarized from unpolarized GWBs than $\Delta_{10}$(PY)5. Although these log Bayes factors do not indicate a strong preference for a PV model, this injection demonstrates that (SR)3(PY)3 can rule out $\Pi=0$ at the $95\%$ confidence level, while $\Delta_{10}$(PY)5 cannot.
In conclusion, these findings highlight that network design strongly impacts parameter constraints, particularly for circular polarization $\Pi$, with (SR)3(PY)3 providing tighter constraints on parity-violating GWBs than $\Delta_{10}$(PY)5.